OpenStudy (anonymous):

Solve the system by the substitution method. xy = 12 x2 + y2 = 40

OpenStudy (anonymous):

Why can't you do your own work

OpenStudy (anonymous):

y dont u pop off

OpenStudy (anonymous):

no thanks

OpenStudy (anonymous):

yea get off my question if u just wanna start trouble toodels

OpenStudy (anonymous):

Thats hilarious sorry! :) But anyways you can use either equation: \[xy = 12 \]\[x= \frac{ 12 }{ y }\] OR \[x^2 + y^2 = 40 \]\[x + y = \sqrt{40}\]\[x = \sqrt{40} - y\] **The y and the x can be switched....

OpenStudy (anonymous):

All you need to do is plug in.

OpenStudy (anonymous):

haha i know .. ok then y does it give me 4 quandrants to pick from ?

OpenStudy (anonymous):

example {( 2, 6), ( 6, 2), ( 2, -6), ( 6, -2)}

OpenStudy (anonymous):

@ilfy214

OpenStudy (anonymous):

From what I solved I got these solutions. 6, 2, -2

OpenStudy (anonymous):

im just confused on plugging them in ..

OpenStudy (anonymous):

Doesn't it depend on how you solved your equation? Also, when I plugged in -2 and 6 for the first problem, I got -12 instead of 12.

OpenStudy (anonymous):

im going w/ b

OpenStudy (anonymous):

I was thinking for along the lines of C?

OpenStudy (anonymous):

i just dont understand how i allways mess that up

OpenStudy (anonymous):

Because in B, if you try using ( 2, -6) and ( -2, 6) you already have a problem for the equation below:\[xy = 12\]\[(2)(-6) = -12\]\[(-2)(6)= -12\]

OpenStudy (anonymous):

Its okay! Just choose the options that give you BOTH positive answers

OpenStudy (anonymous):

thank u i appreciate it !! :)

OpenStudy (anonymous):

You're welcome! :D